# A Short Review for Ontology Learning from Text: Stride from Shallow Learning, Deep Learning to Large Language Models Trend

### v1

@@ Abstract
Ontologies provide formal representation of knowledge shared within Semantic Web applications and Ontology learning from text involves the construction of ontologies from a given corpus of text. In the past years, ontology learning has traversed through shallow learning and deep learning methodologies, each offering distinct advantages and limitations in the quest for knowledge extraction and representation. A new trend of these approaches is relying on large language models to enhance ontology learning. This paper gives a review in approaches and challenges of ontology learning. It analyzes the methodologies and limitations of shallow-learning-based and deep-learning-based techniques for ontology learning, and provides comprehensive knowledge for the frontier work of using large language models to enhance ontology learning. In addition, it proposes several noteworthy future directions for further exploration into the integration of large language models with ontology learning tasks.
@@ 1 Introduction
The purpose of this paper is to give a review in approaches and challenges of ontology learning. It presents the methods and analyzes the limitations of shallow-learning-based and deep-learning-based techniques, and provides comprehensive knowledge for the current work of using large language models to enhance ontology learning. In addition, it proposes several noteworthy future directions for further exploration into the integration of large language models with ontology learning tasks. The rest of this paper is organized as follows: Section 2 defines ontology, ontology learning, and summarises the challenges of ontology learning. Section 3 presents the ontology learning approaches based on shallow learning and deep learning, as well as their limitations. Section 4 presents how large language models contributes to ontology learning procedure recently, and discusses the potential of using large language models to facilitate ontology learning. Section 5 proposes several future directions for further exploration into using large language models to enhance ontology learning. Finally, we conclude in Section 6.
@@ 4.1 Development of Large Language Models
Figure 2: (left) Transformer architecture [85]. (upper right) GPT architecture [86]. (lower right) BERT architecture [89].
@@ 4.2 Concept Extraction
Figure 3: Examples of the template to convert Natural Language sentences into OWL Functional Syntax [6].
[6] fine-tuned a GPT-3 model to convert natural language sentences into OWL Functional Syntax and employed objective and concise examples to fine-tune the model regarding: instances, concepts, as well as class subsumption, domain and range of relations, object properties relationships, disjoint classes, complements, cardinality restrictions. Figure 3 gives some examples of the template. [7] shows that LLMs can leverage their extensive pre-trained knowledge to categorize terms. In particular, LLMs undertake term typing by discerning the contextual placement of a term within a given domain. This can be facilitated through well-crafted prompts in a close or prefix style. A term is presented to the model within a sentence, followed by a [MASK] token indicating the type to be predicted. The model, drawing upon its extensive training, fills in this blank with an appropriate type, effectively classifying the term, with vast data sources. Based on the characteristics of these sources, 8 different prompts have been designed to conduct zero-shot testing on multiple LLMs. In comparison to traditional term typing methods which often rely on manual categorization or simpler NLP techniques, LLMs show the potential in terms of speed, scalability, and contextual accuracy.
@@ 4.3 Relation Extraction
Some studies have achieved good results in the process of ontology learning by using large language models to build taxonomy. [7] evaluated the ability of LLMs to automatically discover the taxonomy of a given knowledge source. 4 top-down taxonomy templates for super-class prediction and 4 bottom-up prompt templates for subclass prediction are defined for LLMs. The results show that LLMs can effectively discover type taxonomies from different knowledge sources. In addition, the performances of LLMs are compared with that of traditional ontology learning methods, such as lexico-syntactic pattern mining and clustering, showing that LLMs outperform these methods in most cases. [100] defined a method for automatically constructing a concept hierarchy for a given domain by querying a large language model as shown in Figure 4. Based on OpenAI’s GPT 3.5, [100] was applied to various domains such as Animals, Drinks, Music, and Plants. The basic idea is to crawl the hierarchy by repeatedly asking the LLM to provide relevant subconcepts of the given concepts that are already in the hierarchy and used an established traversal algorithm to place the new concepts. Further, the LLM was asked to provide a textual description of each concept that they made available for inspection.
Figure 4: Concept Hierarchy Construction [100]

### v2

# A Short Review for Ontology Learning: Stride to Large Language Models Trend

@@ Abstract
Ontologies provide formal representation of knowledge shared within Semantic Web applications. Ontology learning involves the construction of ontologies from a given corpus. In the past years, ontology learning has traversed through shallow learning and deep learning methodologies, each offering distinct advantages and limitations in the quest for knowledge extraction and representation. A new trend of these approaches is relying on large language models (LLMs) to enhance ontology learning. This paper gives a review in approaches and challenges of ontology learning. It analyzes the methodologies and limitations of shallow-learning-based and deep-learning-based techniques for ontology learning, and provides comprehensive knowledge for the frontier work of using LLMs to enhance ontology learning. In addition, it proposes several noteworthy future directions for further exploration into the integration of LLMs with ontology learning tasks.
@@ 1 Introduction
The purpose of this paper is to give a review in approaches and challenges of ontology learning in LLMs era. It presents the methods and analyzes the limitations of shallow-learning-based and deep-learning-based techniques, and provides comprehensive knowledge for the current work of using LLMs to enhance ontology learning. In addition, it proposes several noteworthy future directions for further exploration into the integration of large language models with ontology learning tasks. The rest of this paper is organized as follows: Section 2 defines ontology, ontology learning, and summarises the challenges of ontology learning. Section 3 presents the ontology learning approaches based on shallow learning and deep learning, as well as their limitations. Section 4 presents how large language models contributes to ontology learning procedure recently, and discusses the potential of using large language models to facilitate ontology learning. Section 5 proposes several future directions for further exploration into using large language models to enhance ontology learning. Finally, we conclude in Section 6.
@@ 4.2 Concept Extraction
[6] fine-tuned a GPT-3 model to convert natural language sentences into OWL Functional Syntax and employed objective and concise examples to fine-tune the model regarding: instances, concepts, as well as class subsumption, domain and range of relations, object properties relationships, disjoint classes, complements, cardinality restrictions. [7] shows that LLMs can leverage their extensive pre-trained knowledge to categorize terms. In particular, LLMs undertake term typing by discerning the contextual placement of a term within a given domain. This can be facilitated through well-crafted prompts in a close or prefix style. A term is presented to the model within a sentence, followed by a [MASK] token indicating the type to be predicted. The model, drawing upon its extensive training, fills in this blank with an appropriate type, effectively classifying the term, with vast data sources. Based on the characteristics of these sources, 8 different prompts have been designed to conduct zero-shot testing on multiple LLMs. In comparison to traditional term typing methods which often rely on manual categorization or simpler NLP techniques, LLMs show the potential in terms of speed, scalability, and contextual accuracy.
@@ 4.3 Relation Extraction
Some studies have achieved good results in the process of ontology learning by using large language models to build taxonomy. [7] evaluated the ability of LLMs to automatically discover the taxonomy of a given knowledge source. 4 top-down taxonomy templates for super-class prediction and 4 bottom-up prompt templates for subclass prediction are defined for LLMs. The results show that LLMs can effectively discover type taxonomies from different knowledge sources. In addition, the performances of LLMs are compared with that of traditional ontology learning methods, such as lexico-syntactic pattern mining and clustering, showing that LLMs outperform these methods in most cases. [100] defined a method for automatically constructing a concept hierarchy for a given domain by querying a LLM. Based on OpenAI’s GPT 3.5, [100] was applied to various domains such as Animals, Drinks, Music, and Plants. The basic idea is to crawl the hierarchy by repeatedly asking the LLM to provide relevant subconcepts of the given concepts that are already in the hierarchy and used an established traversal algorithm to place the new concepts. Further, the LLM was asked to provide a textual description of each concept that they made available for inspection.